ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

9. Ordinary Differential Equations and Dynamical Systems

1991 MS Classification: 92 Khrennikov, Andrei, University of Växjö, Sweden The human subconsciousness as a p-adic dynamical system We study a class of dynamical systems over the fields of p-adic numbers that can be used for the description of the process of human thinking. A p-adic dynamical model has been proposed in [1]. This dynamical system is assumed to be located in the human subconsciousness, is controlled by the human consciousness which fixes parameters of the dynamical system in the subconsciousness and transmits to the subconsciousness generating ideas which initiate the iteration of the dynamical system in the subconsciousness. Attractors of this dynamical system are considered by the consciousness as results of the process of thinking. Reference: [1] A. Yu. Khrennikov, Non-Archimedean analysis: quantum paradoxes, dynamical systems and biological models. Kluwer Academic Publishers, Dordrecht, 1997.